# Causally Reliable Concept Bottleneck Models

### v2

@@ 1 Introduction
In recent years, interpretable neural models have become more popular, achieving performance similar to powerful opaque Deep Neural Networks (DNNs) (Alvarez Melis & Jaakkola, 2018; Chen et al., 2019, 2020). Among these, Concept Bottleneck Models (CBMs) (Koh et al., 2020; Zarlenga et al., 2022; Yuksekgonul et al., 2022; Barbiero et al., 2023) guarantee high expressivity and interpretability by enforcing DNNs to reason through a layer of high-level, human-interpretable variables called concepts (e.g., the “color” and “shape” of an object) (Kim et al., 2018; Achtibat et al., 2023; Fel et al., 2023). In CBMs, a neural encoder first maps the raw input to concepts, forming a semantically transparent intermediate representation that is used by a simple decoder for downstream predictions. Beyond transparency, this design allows human experts to intervene on mispredicted concepts at test time to improve downstream task predictions (Espinosa Zarlenga et al., 2024).
However, like standard DNN architectures, CBMs remain pure associative models (Pearl, 2019): their decision-making process reflects statistical correlations within the data rather than real-world causal mechanisms. As a result, they fail to distinguish between spurious correlations and true causal relationships. Recognizing this distinction is fundamental to achieving a reliable scientific understanding, supporting causal reasoning for intervention (Pearl, 2009; Peters et al., 2017), enabling robust generalization under distributional shifts, and the implementation of fairness constraints (Schölkopf et al., 2021; Wang et al., 2022).
To address these limitations, we propose Causally reliable Concept Bottleneck Models (C2BMs): a class of concept-based architectures that enforce reasoning through a “Causal Bottleneck” (Fig. 1) of concepts structured according to a model of the real-world causal mechanisms underlying data generation. C2BMs process information as follows. First, a neural encoder extracts a set of latent representations from raw data. Then, information flows from latent representations through a given causal graph where each node represents an interpretable variable (e.g., “smoker”, “bronchitis”). At inference time, the value of each variable is predicted from its causal parents through an interpretable structural equation, parametrized adaptively by a neural meta-model.
@@ 2 Preliminaries
CBMs (Koh et al., 2020) are interpretable-by-design architectures that explain their predictions using high-level interpretable variables called concepts. Standard CBMs decompose prediction into two stages: a neural encoder maps the input $X$ to a set of intermediate concepts $\mathcal{V}=\{V_{i}\}_{i=1}^{C}$, and a decoder predicts the target $Y$ from $\mathcal{V}$. This yields:
@@
Concept Embedding Models (CEMs) (Zarlenga et al., 2022) enhance CBMs by pairing concepts with high-dimensional embeddings of the form $P(\mathcal{U}\mid\mathcal{V},X)$, where $\mathcal{U}=\{U_{i}\}_{i=1}^{C}$. These embeddings are provided to the decoder to predict the target variable $Y$, enabling the model to achieve performance comparable to standard DNN approaches while maintaining semantic interoperability. Critically, traditional decoders rely on a bipartite structure assumption, wherein all concepts are treated as direct causes of the target, e.g., $Y=f(V_{1},\dots,V_{C})$ for CBMs. This assumption is often overly simplistic for real-world problems. Bringing the reasoning of concept-based architectures closer to real-world mechanisms constitutes the main focus of this work.
@@
A model $\mathcal{M}$ is causally reliable w.r.t. a target phenomenon $T$ if and only if the structure of $\mathcal{M}$’s decision-making process is consistent with the causal mechanisms underlying $T$ (Termine & Primiero, 2024). Although state-of-the-art DNNs and concept-based models offer high expressivity, they lack causal reliability.
@@
The standard framework for modeling causal mechanisms is the structural causal model (SCM) (Bareinboim et al., 2022). An SCM $\mathcal{M}$ is a tuple $\langle\mathcal{V},\,\mathcal{U},\,\mathcal{F},\,P\rangle$, where:
@@
  • =$\mathcal{F}=\{f_{i}\}_{i=1}^{C}$ is a set of functions such that $\>V_{i}=f_{i}(\textsf{PA}_{i},\mathcal{U}_{i})$  for each $i$, where ${\textsf{PA}_{i}\subseteq\mathcal{V}\setminus V_{i}}$, ${\mathcal{U}_{i}\subseteq\mathcal{U}}$, and the entire set $\mathcal{F}$ forms a mapping from $\mathcal{U}$ to $\mathcal{V}$.
@@
Each SCM can be associated with a graphical representation in which nodes correspond to the variables, and edges encode the functional relationships specified by $\mathcal{F}$. Here, we focus on SCMs whose associated graph is a directed acyclic graph (DAG) (Pearl, 1995; Zaffalon et al., 2020b). In most cases, the underlying DAG is unknown and must be inferred from observational data, a process known as causal discovery (Peters et al., 2017; Zanga et al., 2022). However, methods based solely on observational data cannot generally guarantee the identification of a unique DAG (Peters et al., 2017). The set of candidate DAGs can be refined by incorporating additional information, which we refer to as background knowledge (Andrews et al., 2020; Abdulaal et al., 2023). This can be drawn from a range of sources, such as human experts, structured repositories of information (e.g., domain ontologies), or “unstructured" samples of information (e.g., scientific papers or other documentation).
@@ 3 Related works
Traditional concept-based architectures impose a strict bipartite structure in which concept neuron activations are assumed to directly cause task outputs (Koh et al., 2020; Yuksekgonul et al., 2022; Kim et al., 2023; Oikarinen et al., 2023; Yang et al., 2023; Barbiero et al., 2023; Vandenhirtz et al., 2024). This strong, often unrealistic assumption can lead to misleading explanations. For example, attributing a lung cancer diagnosis to both a ‘cough’ and ‘smoker’ concept could risk the false interpretation that reducing coughing could reduce cancer risk. Moreover, most CBMs assume independence among concepts, which is unrealistic, as it ignores natural co-occurrences (e.g., ‘smoke’ and ‘fire’) and prevents improvements in one concept from propagating to related concepts during interventions. Stochastic CBM (SCBM) (Vandenhirtz et al., 2024) attempts to relax this assumption by learning a joint distribution over concepts. However, this approach captures only associations rather than causal relations, making it vulnerable to spurious correlations in the data. To date, no methodology exists for structuring the concept bottleneck according to a reliable causal model.
Recent approaches like DiConStruct (Moreira et al., 2024), aim to improve this aspect by generating causal graphs linking concepts to opaque DNN predictions. However, DiConStruct is a post-hoc method that may misalign with the original DNN’s outputs and relies solely on observational data, neglecting background knowledge and resulting in under-determined causal structures. Other architectures, such as Neural Causal Models (Ke et al., 2019) and Neural Causal Abstractions (Xia & Bareinboim, 2024), impose even stronger assumptions, requiring access to either the true causal graph or a low-resolution structural causal model, which are impractical in many cases.
@@ 4.1 Causally reliable Concept Bottleneck Models
  • $\mathcal{M}_{\bm{\Theta}}$ is a parametric SCM $\langle\mathcal{V},\mathcal{U},\mathcal{F}_{\bm{\Theta}},P(\mathcal{U}|X)\rangle$ (see Sec. 2), where we assume a parametric form for the functions’ set. Specifically, the structure of the functions is determined by the connectivity of $\mathbf{G}$, and the parameters $\bm{\Theta}$ are predicted from $\mathcal{U}$ by a neural meta-model.
@@ 4.2 Model instantiation
Our pipeline (see Fig. 2) addresses the following sub-problems: (i) causal graph construction, which includes concept discovery, concept labeling (Sec. 4.2.1), and causal graph discovery (Sec. 4.2.2); and (ii) training of the neural parameters of the encoder and the meta-model determining the structural equations (Sec. 4.3). In the following, we outline our implementation for each sub-problem.
@@
Concepts and causal graph constitute an input for C2BM, which could remain agnostic to how they are obtained, e.g., provided by human experts. Notably, alternative or novel approaches may be employed, provided they solve the same problems.
@@ 4.2.1 Concept discovery and labeling
In the CBM community, automated concept discovery and labeling using Large Language Models (LLMs) has become a standard solution when human supervision is unavailable (Oikarinen et al., 2023; Yang et al., 2023; Srivastava et al., 2024; Yamaguchi & Nishida, 2025). In our implementation, we follow the label-free CBM approach from Oikarinen et al. (2023), where concepts are discovered by querying an LLM for those most relevant to the task. We then apply a filtering procedure to retain only concepts that meet criteria such as brevity, distinctiveness (i.e., not too similar to each other or the target), and presence in the training data.
Once $\mathcal{V}$ are selected, we label the dataset with variable annotations $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{v}_{i})\}_{i=1}^{N}$ to supervise concept learning. To do so, we adopt a strategy similar to Oikarinen et al. (2023), leveraging a pre-trained contrastive vision-language model, such as CLIP (Radford et al., 2021). This projects both data samples and the discovered concept names into a shared embedding space and computes their alignment to generate concept labels. Full implementation details are provided in App. B.
@@ 4.2.2 Causal graph discovery
Let $\mathbf{G}^{*}$ be the true, unknown, graph over a set of variables $\mathcal{V}$ from which a dataset $\mathcal{D}$ was generated. The causal discovery problem consists in recovering $\mathbf{G}^{*}$ from the observed dataset $\mathcal{D}$ (Zanga et al., 2022).
As anticipated in Sec. 2, a promising direction for addressing this problem is to combine standard causal discovery algorithms with knowledge-base querying. In our pipeline, we focus on a well-known class of methods that recover an equivalence class of graphs from data, referred to as the Markov equivalence class (MEC) (Spirtes et al., 2001; Pearl, 2009; Zanga et al., 2022). This class can be compactly represented as a Completed Partially Directed Acyclic Graph (CPDAG), an extension of a DAG where edges remain unoriented when there is insufficient evidence in the data to infer causal direction. This is a desirable property as it prevents incorrect (spurious) orientations based on the data alone. Specifically, we apply the Greedy Equivalence Search (GES) (Chickering, 2002) algorithm, which we found performed well empirically (see App. G.4). Then, we leverage a pre-trained LLM to assess each undirected edge in the CPDAG, orienting the ones corresponding to true causal relationships while discarding those originating from spurious correlations. To improve the robustness and generalizability of this approach, we pair the LLM with a Retrieval Augmented Generation (RAG) technique, which is known to reduce hallucinations and can provide problem-specific knowledge. To further improve robustness, we repeat each query 10 times selecting the most frequent outcome. Implementation details are provided in App. C.
@@ 4.3 Structural equations and model training
We model the structural functions $f_{i}\in\mathcal{F}_{\bm{\Theta}}$ describing the causal mechanisms relating each endogenous variable to its endogenous parents as weighted linear sums, i.e., for each $i$:
@@
where ${\textsf{PA}_{i}}$ denotes the set of endogenous parents of $V_{i}$. For the parameters ${\bm{\theta}_{f_{i}}}$, we do not learn a single parameterization; instead, these are adaptively inferred for each different realization of $X$, by a neural meta-model $\mathbf{r}(\cdot)$, based on the values of the exogenous variables $\mathcal{U}$ and the graph connectivity. In our implementation, we consider the meta-model to be composed of separate meta-models $\mathbf{r}_{i}(\cdot)$ (e.g., independent DNNs), each taking as input a separate exogenous variable:
@@
The design in Eq. 2-3 improves both interpretability and expressivity. To aid (mechanistic) interpretability, structural equations take a linear form (Eq. 2): the value of each is a weighted linear combination of its parents. The adaptive re-parameterization of the equation’s weights performed by the meta-model $\mathbf{r}(\cdot)$ allows the model to also approximate non-linear relationships among endogenous variables (see App. D, which also includes a proof that C2BM is a universal approximator, regardless of the underlying causal graph). This idea is in line with existing literature on interpretability (Ribeiro et al., 2016; Alvarez Melis & Jaakkola, 2018) and concept-based methods (Barbiero et al., 2023)  [^1] This idea also aligns with Zaffalon et al. (2020a); Balke & Pearl (1994), where exogenous variables are used to represent relationships between endogenous variables when the structural equations are unknown.
@@
The training of C2BM consists of learning the neural parameters of the encoder $\mathbf{g}(\cdot)$ and meta-model $\mathbf{r}(\cdot)$ end-to-end from the input data. We formalize this by modeling the joint conditional distribution $P(\mathcal{V},\mathcal{U},\bm{\Theta}\mid X,\mathcal{E})$ which factorizes as:
$P(\mathcal{V},\mathcal{U},\bm{\Theta}\mid X,\mathcal{E})=\overbrace{P(\mathcal{V}\mid\mathcal{U};\bm{\Theta})}^{\text{endogenous}}\,\overbrace{P(\bm{\Theta}\mid\mathcal{E},\mathcal{U})}^{\text{structural equation}}\,\overbrace{P(\mathcal{U}\mid X)}^{\text{exogenous}}$ (4)
where: $P(\mathcal{U}\mid X)$ represents the exogenous encoder $\mathbf{g}(\cdot)$; $P(\bm{\Theta}\mid\mathcal{E},\mathcal{U})$ represents the neural meta-model $\mathbf{r}(\cdot)$ predicting the structural equations’ parameters using the given causal connections and the exogenous variables; $P(\mathcal{V}\mid\mathcal{U};\bm{\Theta})$ represents a causally reliable classifier leveraging the structural equations $\mathcal{F}_{\bm{\Theta}}$ to predict the values of the endogenous variables. Under the Markov condition imposed by the C2BM causal graph, the causally-reliable classifier can be re-written as a product of independent distributions i.e.,
@@
We clarify that we do not claim to identify the true structural functions. Instead, C2BM numerically approximates the outcomes as if they were generated by the underlying (unknown) structural equations. This approximation, along with C2BM’s DAG, is sufficient to compute reliable interventions, which is a core objective in the concept-based community (Poeta et al., 2023).

### v3

@@ 1 Introduction
In recent years, interpretable neural models have become more popular, achieving performance similar to powerful opaque Deep Neural Networks (DNNs) (alvarez2018towards; chen2019looks; chen2020concept). Among these, Concept Bottleneck Models (CBMs) (koh2020; zarlenga2022concept; yuksekgonul2022post; barbiero2023interpretable) guarantee high expressivity and interpretability by enforcing DNNs to reason through a layer of high-level, human-interpretable variables called concepts (e.g., the “color” and “shape” of an object) (kim2018interpretability; achtibat2023attribution; fel2023craft). In CBMs, a neural encoder first maps the raw input to concepts, forming a semantically transparent intermediate representation that is used by a simple decoder for downstream predictions. Beyond transparency, this design allows human experts to intervene on mispredicted concepts at test time to improve downstream task predictions (espinosa2024learning).
However, like standard DNN architectures, CBMs remain pure associative models (Pearl2019): their decision-making process reflects statistical correlations within the data rather than real-world causal mechanisms. As a result, they fail to distinguish between spurious correlations and true causal relationships. Recognizing this distinction is fundamental to achieving a reliable scientific understanding, supporting causal reasoning for intervention (Pearl2009; peters2017elements), enabling robust generalization under distributional shifts, and the implementation of fairness constraints (scholkopf2021toward; wang2022out).
To address these limitations, we propose Causally reliable Concept Bottleneck Models (C2BMs): a class of concept-based architectures that enforce reasoning through a “Causal Bottleneck” (Fig. 1) of concepts structured according to a model of the real-world causal mechanisms underlying data generation. C2BMs process information as follows. First, a neural encoder extracts a set of latent representations from raw data. Then, information flows from latent representations through a given causal graph where each node represents an interpretable variable (e.g., “smoker”, “bronchitis”). At inference time, the value of each variable is predicted from its causal parents through an interpretable structural equation, parametrized adaptively by a hypernetwork.
@@ 2 Preliminaries
CBMs (koh2020) are interpretable-by-design architectures that explain their predictions using high-level interpretable variables called concepts. Standard CBMs decompose prediction into two stages: a neural encoder maps the input $X$ to a set of intermediate concepts $\mathcal{V}=\{V_{i}\}_{i=1}^{C}$, and a decoder predicts the target $Y$ from $\mathcal{V}$. This yields:
@@
Concept Embedding Models (CEMs) (zarlenga2022concept) enhance CBMs by pairing concepts with high-dimensional embeddings of the form $P(\mathcal{U}\mid\mathcal{V},X)$, where $\mathcal{U}=\{U_{i}\}_{i=1}^{C}$. These embeddings are provided to the decoder to predict the target variable $Y$, enabling the model to achieve performance comparable to standard DNN approaches while maintaining semantic interpretability. Critically, traditional decoders rely on a bipartite structure assumption, wherein all concepts are treated as direct causes of the target, e.g., $Y=f(V_{1},\dots,V_{C})$ for CBMs. This assumption is often overly simplistic for real-world problems. Bringing the reasoning of concept-based architectures closer to real-world mechanisms constitutes the main focus of this work.
@@
A model $\mathcal{M}$ is causally reliable w.r.t. a target phenomenon $T$ if and only if the structure of $\mathcal{M}$’s decision-making process is consistent with the causal mechanisms underlying $T$ (Termine2023Causal). Although state-of-the-art DNNs and concept-based models offer high expressivity, they lack causal reliability.
@@
The standard framework for modeling causal mechanisms is the structural causal model (SCM) (bareinboim2022pearl). An SCM $\mathcal{M}$ is a tuple $\langle\mathcal{V},\,\mathcal{U},\,\mathcal{F},\,P\rangle$, where:
@@
  • $\mathcal{F}=\{f_{i}\}_{i=1}^{C}$ is a set of functions such that
    $V_{i}=f_{i}(\textsf{PA}_{i},\,\mathcal{U}_{i})\quad\forall i=1,\dots,C$ (2)
    where $\textsf{PA}_{i}\subseteq\mathcal{V}\setminus V_{i}$} is the set of the endogenous parents of $V_{i}$, $\mathcal{U}_{i}\subseteq\mathcal{U}$} is an exogenous parent summarizing all the information influencing $V_{i}$ that is not explicitly represented in $\mathcal{V}$, and the entire set $\mathcal{F}$ forms a mapping from $\mathcal{U}$ to $\mathcal{V}$.
@@
Each SCM can be associated with a graphical representation in which nodes correspond to the variables, and edges encode the functional relationships specified by $\mathcal{F}$. Here, we focus on SCMs whose associated graph is a directed acyclic graph (DAG) (pearl1995causal; zaffalon2020structural). In most cases, the underlying DAG is unknown and must be inferred from observational data, a process known as causal discovery (peters2017elements; zanga2022). However, methods based solely on observational data cannot generally guarantee the identification of a unique DAG (peters2017elements). The set of candidate DAGs can be refined by incorporating additional information, which we refer to as background knowledge (andrews2020completeness; abdulaal2023causal). This can be drawn from a range of sources, such as human experts, structured repositories of information (e.g., domain ontologies), or “unstructured” samples of information (e.g., scientific papers or other documentation).
@@ 3 Related works
Traditional concept-based architectures impose a strict bipartite structure in which concept neuron activations are assumed to directly cause task outputs (koh2020; yuksekgonul2022post; kim2023probabilistic; oikarinen2023; yang2023language; barbiero2023interpretable; vandenhirtz2024stochastic). This strong, often unrealistic assumption can lead to misleading explanations. For example, attributing a lung cancer diagnosis to both a ‘cough’ and ‘smoker’ concept could risk the false interpretation that reducing coughing could reduce cancer risk. Moreover, most CBMs assume independence among concepts, which is unrealistic, as it ignores natural co-occurrences (e.g., ‘smoke’ and ‘fire’) and prevents improvements in one concept from propagating to related concepts during interventions. Stochastic CBM (SCBM) (vandenhirtz2024stochastic) and Concept Graph Models (dominici2025causal) attempt to relax this assumption. However, these approaches capture only associations rather than causal relations, making it vulnerable to spurious correlations in the data. To date, no methodology exists for structuring the concept bottleneck according to a reliable causal model.
Recent approaches like DiConStruct (moreira2024diconstruct), aim to improve this aspect by generating causal graphs linking concepts to opaque DNN predictions. However, DiConStruct is a post-hoc method that may misalign with the original DNN’s outputs and relies solely on observational data, neglecting background knowledge and resulting in under-determined causal structures. Other architectures, such as Neural Causal Models (ke2019learning) and Neural Causal Abstractions (xia2024neural), impose even stronger assumptions, requiring access to either the true causal graph or a low-resolution structural causal model, which are impractical in many cases.
@@ 4.1 Causally reliable Concept Bottleneck Models
  • $\mathcal{M}_{\bm{\Theta}}$ is a parametric SCM $\langle\mathcal{V},\mathcal{U},\mathcal{F}_{\bm{\Theta}},P(\mathcal{U}|X)\rangle$ (see Sec. 2), where we assume a parametric form for the functions’ set. Specifically, the structure of the functions is determined by the connectivity of $\mathbf{G}$, and the parameters $\bm{\Theta}$ are predicted from $\mathcal{U}$ by a hypernetwork.
@@ 4.2 Model instantiation
Our pipeline (see Fig. 2) addresses the following sub-problems: (i) causal graph construction, which includes concept discovery, concept labeling (Sec. 4.2.1), and causal graph discovery (Sec. 4.2.2); and (ii) training of the neural parameters of the encoder and the hypernetwork determining the structural equations (Sec. 4.3). In the following, we outline our implementation for each sub-problem. Specifically, building C2BM’s individual prerequisites will be mostly based on prior work. Note that integrating them into a coherent, automated pipeline is instead part of this paper’s contributions.
@@
Concepts and causal graph constitute an input for C2BM, which could remain agnostic to how they are obtained, e.g., provided by human experts. Notably, alternative or novel approaches may be employed, provided they solve the same problems (loula2025syntactic).
@@ 4.2.1 Concept discovery and labeling
In the CBM community, automated concept discovery and labeling using Large Language Models (LLMs) has become a standard solution when human supervision is unavailable (oikarinen2023; yang2023language; srivastava2024vlg; yamaguchi2025explanation). In our implementation, we follow the label-free CBM approach from oikarinen2023, where concepts are discovered by querying an LLM for those most relevant to the task. We then apply a filtering procedure to retain only concepts that meet criteria such as brevity, distinctiveness (i.e., not too similar to each other or the target), and presence in the training data.
Once $\mathcal{V}$ are selected, we label the dataset with variable annotations $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{v}_{i})\}_{i=1}^{N}$ to supervise concept learning. To do so, we adopt a strategy similar to oikarinen2023, leveraging a pre-trained contrastive vision-language model, such as CLIP (radford2021learning). This projects both data samples and the discovered concept names into a shared embedding space and computes their alignment to generate concept labels. Full implementation details are provided in App. B.
@@ 4.2.2 Causal graph discovery
Let $\mathbf{G}^{*}$ be the true, unknown, graph over a set of variables $\mathcal{V}$ from which a dataset $\mathcal{D}$ was generated. The causal discovery problem consists in recovering $\mathbf{G}^{*}$ from the observed dataset $\mathcal{D}$ (zanga2022).
As anticipated in Sec. 2, a promising direction for addressing this problem is to combine standard causal discovery algorithms with knowledge-base querying. In our pipeline, we focus on a well-known class of methods that recover an equivalence class of graphs from data, referred to as the Markov equivalence class (MEC) (Spirtes2001; Pearl2009; zanga2022).This class can be compactly represented as a Completed Partially Directed Acyclic Graph (CPDAG), an extension of a DAG where edges remain unoriented when there is insufficient evidence in the data to infer causal direction. This is a desirable property as it prevents incorrect (spurious) orientations based on the data alone. Specifically, we apply the Greedy Equivalence Search (GES) (chickering2002) algorithm, which we found performed well empirically (see App. G.3). Then, we leverage a pre-trained LLM to assess each undirected edge in the CPDAG, orienting the ones corresponding to true causal relationships while discarding those originating from spurious correlations. To improve the robustness and generalizability of this approach, we pair the LLM with a Retrieval Augmented Generation (RAG) technique, which is known to reduce hallucinations and can provide problem-specific knowledge. To further improve robustness and performance, we repeat each query 10 times, selecting the most frequent outcome (wang2022self). Implementation details are provided in App. C.
@@ 4.3 Structural equations and model training
We model the structural functions $f_{i}\in\mathcal{F}_{\bm{\Theta}}$ describing the causal mechanisms relating each endogenous variable  [^1] Root variables are predicted from the exogenous variables $\mathcal{U}$ using a neural network. Further details are provided in App. D. to its endogenous parents as weighted linear sums, i.e., for each $i$:
@@
where $\textsf{PA}_{i}$} denotes the set of endogenous parents of $V_{i}$ [^2] We refer to $V_{i}$ as a variable to allow for a more general formalization. In a classification setting, concepts assume categorical values; hence $V_{i}$ represents the probability of a concept state activation.. For the parameters $\bm{\theta}_{f_{i}}$}, we do not learn a single parameterization; instead, these are adaptively inferred for each different realization of $X$, by a hypernetwork $\mathbf{r}(\cdot)$ (ha2017hypernetworks; barbiero2023interpretable; debot2024interpretable), based on the values of the exogenous variables $\mathcal{U}$ and the graph connectivity. In our implementation, we consider separate hypernetworks $\mathbf{r}_{i}(\cdot)$ (e.g., independent DNNs), each taking as input a separate exogenous variable:
@@
The design in Eq. 3 and 4 improves both interpretability and expressivity. To aid (mechanistic) interpretability, structural equations take a linear form (Eq. 3): the value of each is a weighted linear combination of its parents. The adaptive re-parameterization of the equation’s weights performed by the hypernetwork $\mathbf{r}(\cdot)$ allows the model to also approximate non-linear relationships among endogenous variables (see App. D, which also includes a proof that C2BM is a universal approximator, regardless of the underlying causal graph). This idea is in line with existing literature on interpretability (ribeiro2016should; alvarez2018towards) and concept-based methods (barbiero2023interpretable)  [^3] This idea also aligns with balke1994counterfactual and zaffalon2020, where exogenous variables are used to represent relationships between endogenous variables when the structural equations are unknown.
@@
The training of C2BM consists of learning the neural parameters of the encoder $\mathbf{g}(\cdot)$ and hypernetwork $\mathbf{r}(\cdot)$ end-to-end from the input data. We formalize this by modeling the joint conditional distribution $P(\mathcal{V},\mathcal{U},\bm{\Theta}\mid X,\mathcal{E})$ which factorizes as:
[…]
where: $P(\mathcal{U}\mid X)$ represents the exogenous encoder $\mathbf{g}(\cdot)$; $P(\bm{\Theta}\mid\mathcal{E},\mathcal{U})$ represents the hypernetwork $\mathbf{r}(\cdot)$ predicting the structural equations’ parameters using the given causal connections and the exogenous variables; $P(\mathcal{V}\mid\mathcal{U};\bm{\Theta})$ represents a causally reliable classifier leveraging the structural equations $\mathcal{F}_{\bm{\Theta}}$ to predict the values of the endogenous variables. Under the Markov condition imposed by the C2BM causal graph, the causally-reliable classifier can be re-written as a product of independent distributions i.e.,
@@
We clarify that we do not claim to identify the true structural functions. Instead, C2BM numerically approximates the outcomes as if they were generated by the underlying (unknown) structural equations. This approximation, along with C2BM’s DAG, is sufficient to compute reliable interventions, which is a core objective in the concept-based community (Poeta2023; steinmann2024learning).